Dividends Per Share

Seacrest Company has 20,000 shares of cumulative preferred 2% stock, $150 par and 50,000 shares of $25 par common stock. The following amounts were distributed as dividends:

Year 1 $150,000
Year 2 48,000
Year 3 180,000

Determine the dividends per share for preferred and common stock for each year. Round all answers to two decimal places. If an answer is zero, enter '0'.

Year 1 Year 2 Year 3
Preferred stock (Dividends per share) $ $ $
Common stock (Dividends per share) $ $ $

Entries for Cash Dividends

The declaration, record, and payment dates in connection with a cash dividend of $40,000 on a corporation's common stock are February 1, March 18, and May 1.

Journalize the entries required on each date. If no entry is required, select "No Entry Required" and leave the amount boxes blank.

Feb. 1 Cash Dividends
Cash Dividends Payable
Mar. 18
May. 1
Answer #1

1. DIVIDEND PER SHARE.

2% Preferred shares = 20,000 x 150

2% preferred shares = $ 3,000,000.

Common shares = 50,000 x 25

Common shares = $ 1,250,000.

Cumulative preference shares are those shares which entitled to get any arrear dividend and current dividend amount first before giving to common shareholders. If any amount left after giving arrear and current dividend amount to them, balance any is given to common shares.

Fixed dividend to preferred shares = 20,000 x 150 x 2%

Fixed dividend to preferred shares = $ 60,000 per year.

In year 1,

Dividend paid = $ 150,000.

Dividend to preferred shares = $ 60,000.

Dividend per share = dividend amount/number of shares.

Dividend per share = 60,000/20,000

Dividend per share = $ 3 per share.

Dividend to common shares = $ 90,000.

Dividend per share = 90,000/50,000

Dividend per share = $ 1.80 per share.

In year 2,

Dividend paid = $ 48,000.

Dividend to preferred shares = $ 48,000.

Dividend per share = 48,000/20,000

Dividend per share = $ 2.40 per share.

Dividend to common shares = $ 0.

Dividend per share = $ 0.

Here, preferred shares have right to get fixed annual dividend of $ 60,000 but only $ 48,000 is paid. So this full is given to them and balance $ 12,000 will be paid in next year.

In year 3,

Dividend paid = $ 180,000.

Dividend paid to preference shares = 12,000+60,000

Dividend to preference shares = $ 72,000

Dividend per share = 72,000/20,000

Dividend per share = $ 3.60 per share.

Dividend to common shares = $ 108,000.

Dividend per share = 108,000/50,000

Dividend per share = $ 2.16 per share.

2. ENTRIES FOR CASH DIVIDENDS.

Declaration date = February 1

Record date = March 18

Payment date = May 1

Dividend amount = $ 40,000.

Date Accounts Debit Credit
Feb 1 Retained earnings $ 40,000
To Cash dividend payable $ 40,000
( declaration of dividend)
March 18 NO ENTRY NO ENTRY NO ENTRY
( Record of date)
May 1 Cash dividend payable $ 40,000
To Cash $ 40,000
( payment of dividend)

SUMMARY:

1. For year 1.

Dividend per share of preferred shares = $ 3 per share.

Dividend per share of common shares = $ 1.80 per share.

For year 2.

Dividend per share of preferred shares = $ 2.40 per share.

Dividend per share of common shares = $ 0 per share.

For year 3.

Dividend per share of preferred shares = $ 3.60 per share.

Dividend per share of common shares = $ 2.16 per share.

2. On declaration date, retained earnings account is debited and dividend payable is credited. On record date no entry is to be passed. On payment date the liability of dividend payable is closed by debiting and cash is credited.